a rectangular parking lot has a length that is 3 yards greater than the width. the ares of the parking lot is 180 square yards. find the length and width.
Answer by rapaljer(4671)

Let x= width
x+3 = length

Area = W*L= 180
x(x+3) = 180

This will be quadratic, so set it equal to zero:
x^2 + 3x -180=0

Hopefully, it FACTORS!!
(x+15)(x-12)=0
x=-15 or x=12

Reject the negative answer, which leaves

x=12 yards width
x+3= 15 yards length

Check: Area= 12*15= 180 square yards!

If you need help either solving Quadratic Equations or Factoring Trinomials, please see my own website for FREE non-traditional explanations of these very important topics. Do a "Bing" or "Google" search for my last name "Rapalje". Near the top of the search list, find "Rapalje Homepage", and near the top of my Homepage, look for the link "Basic, Intermediate, and College Algebra: One Step at a Time." Choose "Basic Algebra", and look in "Chapter 2" for the appropriate topic that you need extra help with. My own students ALL thought my explanations were a LOT easier than the traditional explanations.

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus